(2x+5)(x2-2)=0


1
Expert's answer
2021-06-29T17:39:07-0400

(2x+5)(x2-2)=0

Steps:

  1. (2x+5)=0
  2. 2x=-5
  3. x1= "-\frac{5}{2}"
  4. (x2-2)=0
  5. x2=2
  6. x2= "\sqrt{2}"
  7. x3= "-\sqrt{2}"


The solutions are:

x1= "-\frac{5}{2}" , x2= "\sqrt{2}" , x3= "-\sqrt{2}"




Plotting: